Using the 6 ordinate scheme analyze harmonically the data to third harmonics
X 0 π/3 2π/3 π 4π/3 5π/3 2π
Y 10 12 15. 20. 17. 11. 10
1
Expert's answer
2021-07-19T16:23:22-0400

"\sum_{k=1}^7y_k=95"

"\sum_{k=1}^7y_kcos(x_k)=-4.5"

"\sum_{k=1}^7y_ksin(x_k)=-0.866"

"\sum_{k=1}^7y_kcoc(2x_k)=12.5"

"\sum_{k=1}^7y_ksin(2x_k)=2.598"

we can get the values of "a_i" and "b_i" as follows:

"a_0={1\over 7}\times95=13.57"

"a_1={2\over 7}\times(-4.5)=-1.29"

"a_2={2\over 7}\times 12.5=3.57"

"b_1={2\over 7}\times(-0.866)=-0.25"

"b_2={2\over 7}\times 2.598=0.74"

Substituting those values into Fourier series

"f(x)=a_0+\sum_{n=1}^\infin(a_ncos(nx)+b_nsin(nx))" gives

"f(x)=a_0+a_1cos(x)+a_2cos(2x)+b_1sin(x)+b_2sin(2x)"

"\therefore y=13.57-1.29cos(x)+3.75cos(2x)-0.25sin(x)+0.74sin(2x)"